OpenStudy (anonymous):

8x + 2y= 10 Solve the equation for y

OpenStudy (karatechopper):

ok so get 8x to the other side what equation do u get

OpenStudy (karatechopper):

fine the answer is y=-4x+5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

For others who come along and don't know how to solve this question...it's asking you to solve for y. Therefore, you have to make everything else equal y. 1. 8x + 2y = 10 2. Move 8x over to the other side: 2y = 10 - 8x 3. Divide by 2 to solve for y: y=(10-8x)/2 --> y = 5 - 4x. That's all. It looks complicated because you have both x and y, but it really doesn't matter what x is...just what y is.